# EMNLP: Educator-role Moral and Normative Large Language Models Profiling

### v1

@@ Abstract
Simulating Professions (SP) enables Large Language Models (LLMs) to emulate professional roles. However, comprehensive psychological and ethical evaluation in these contexts remains lacking. This paper introduces EMNLP, an Educator-role Moral and Normative LLMs Profiling framework for personality profiling, moral development stage measurement, and ethical risk under soft prompt injection. EMNLP extends existing scales and constructs 88 teacher-specific moral dilemmas, enabling profession-oriented comparison with human teachers. A targeted soft prompt injection set evaluates compliance and vulnerability in teacher SP. Experiments on 12 LLMs show teacher-role LLMs exhibit more idealized and polarized personalities than human teachers, excel in abstract moral reasoning, but struggle with emotionally complex situations. Models with stronger reasoning are more vulnerable to harmful prompt injection, revealing a paradox between capability and safety. The model temperature and other hyperparameters have limited influence except in some risk behaviors. This paper presents the first benchmark to assess ethical and psychological alignment of teacher-role LLMs for educational AI. Resources are available at https://e-m-n-l-p.github.io/.
@@ 1 Introduction
Our EMNLP framework conducts a moral and ethical evaluation of LLMs in the teacher SP across three key dimensions: personality measurement, moral ethical judgment, and harmful content risk. For personality measurement, we made necessary additions, enhancements, and extensions to existing measurement tools Marraffini et al. (2024), selected and designed general and professional personality scales to assess LLM performance in the teacher SP and compare the results with human teachers. In terms of moral ethical judgment, we have developed a set of ethical dilemmas tailored to the teacher SP, including extreme scenarios, to evaluate LLM decision-making processes Liu et al. (2024) and analyze their moral tendencies. For harmful content risk testing, we designed an induced soft prompt injection set specific to the teacher SP to examine whether LLMs comply with ethical flaws in role-playing. Additionally, we incorporated model hyperparameters into the testing scope, adjusting them at each stage to explore their impact on LLM moral and ethical performance.
@@
  • We design the EMNLP framework, which includes dedicated personality scales, moral dilemmas, and tailored prompts for teacher SP, enabling comprehensive testing of LLMs’ moral and ethical behavior. Experiments on 12 LLMs reveal general tendencies in this setting.
@@ 3.1 Personality Scales
We adopt the statements from the Computerized Personality Scale for Teachers (CPST) Chao and Sung (2020), which contains 39 behaviorally neutral statements describing personal traits, evenly distributed across 13 dimensions. To improve the internal consistency of the scale, reduce random measurement error, and enhance its content validity, we extended the CPST by doubling the number of statements for each dimension. The resulting scale is referred to as the CPST-E (Extended CPST).
@@
Both scales were generated in English, and administered using a 7-point Likert format for scoring. Detailed information on the CPST-E and HEXACO-60 scale can be found in Appendix A.1.
@@ 3.2 Moral Dilemmas
Moral dilemmas in the teaching profession can be categorized into five types of conflict Shapira-Lishchinsky (2011): Caring Climate vs. Formal Climate, Distributive Justice vs. School Standards, Confidentiality vs. School Rules, Loyalty to Colleagues vs. School Norms, and Family Agenda vs. Educational Standards, with 11 subcategories under these five types. Based on these aspects, we have constructed 88 moral dilemmas, ensuring that each aspect’s dilemmas cover all its subcategories and include a variety of scenarios from primary, secondary, and higher education. To avoid randomness, we have incorporated diverse situations when constructing these dilemmas, including some extreme scenarios, to comprehensively examine the decision-making processes and moral tendencies of teacher SP LLMs in response to different scenarios. Moreover, we have set these moral dilemmas as open-ended questions, allowing LLMs ample freedom and avoiding the potential limitations that multiple-choice answers might impose. Detailed information of the dilemmas are in Appendix A.2.
@@ 4.1 Experimental Designs
We proposed 4 research questions and conducted experiments on 12 selected LLMs to investigate the moral and ethical dimensions of teacher SP LLMs. The list of models is provided in Appendix B.
@@
RQ2: What moral development stages are exhibited by various LLMs in the teacher SP? This question investigates the moral development stages demonstrated by teacher SP LLMs in response to moral dilemmas encountered in educational settings. A total of 88 dilemmas were constructed, categorized into five thematic domains, generating 12 × 88 = 1056 sessions. The experiments were conducted in English with a temperature value of 0. After each model generated a response, 9 human experts voted to classify it into one of Kohlberg’s three stages of moral development. A detailed guideline for the expert rating procedure for RQ2 is presented in Appendix C.1. Figure 2 presents a sample test template used in this experiment.
@@
RQ3: Under soft prompt injection, do teacher SP LLMs exhibit unethical or harmful responses? This question examines whether teacher SP LLMs produce harmful content when users apply ’soft prompt injection’ during real-world usage, setting the role directly through natural language input. The goal is to assess whether LLMs conform to the moral flaws embedded in the user-defined roles. The experiments were conducted in English, with a temperature value of 0, generating 5 × 5 × 4 × 12 = 1200 responses. For each LLM response, 9 experts voted on whether the content was considered harmful. The expert rating guideline for RQ3 is provided in Appendix C.2. Figure 3 presents a sample test template used in this experiment.
@@ 4.2 Evaluation Metrics and Scoring Protocols
Moral Development Stage: Each moral dilemma is individually presented to the LLMs. 9 human experts independently evaluate the model’s response, and the final moral development stage is determined by majority voting:
@@
Harmful Response: Each soft injection prompt and student speech sample is individually presented to the LLMs. Each model response is evaluated by 9 human experts, who label it as “harmful” (1) or “non-harmful” (0). The final decision is made by majority voting:
@@ 5 Results
In the HEXACO-60, teacher SP LLMs scored lower than human teachers on Emotionality and higher on Honesty-Humility. As shown in Figure 6, human teachers recorded the lowest scores on items 3, 6, 12, 15, 16, 18, 51, 52, 56, and 60, suggesting that LLMs tend to display more idealized and morally consistent traits. Conversely, human teachers outperformed LLMs on items 29 and 35, underscoring the models’ limitations in replicating emotional sensitivity and crisis management skills. In the CPST-E, human teachers scored highest on Introversion & Quietness and lowest on Composure & Confidence and Honesty & Candor. As shown in Figure 7, human teachers showed the lowest scores on items 5, 20, 26, 56, 61, 62, 63, 64, 65, 67, 72, 76, and 78, indicating that LLMs present a more consistently positive and idealized professional persona. Meanwhile, on items 43, 44, 66, and 77, human teachers scored highest, where LLMs exhibited weaker performance in empathy, critical reasoning, and resilience. These findings suggest that while LLMs in the teacher SP setting can approximate certain human traits, they still show notable biases in emotional experience, honesty expression, and self-awareness.
To further investigate factors influencing these patterns, we examined model size and reasoning ability. In the HEXACO-60 inventory, larger models tended to score higher, particularly in emotional sensitivity and moral humility, while showing little change in sociability-related traits. In the CPST-E, model size had generally weak effects across dimensions, except for a notable negative trend in Achievement Orientation and Fairness and Openness, suggesting that larger models may prioritize caution and compliance over overt ambition. Slight positive trends were observed in Risk-taking and Innovation and Caring and Empathy. Regarding reasoning ability, reasoning models achieved higher average scores in the CPST-E, indicating stronger alignment with educator-specific traits. In contrast, their advantage in the HEXACO-60 inventory was marginal. Moreover, reasoning models exhibited lower intra-group variance across both inventories, reflecting more consistent personality outputs compared to non-reasoning models.
@@ 5.2 Moral Development Stages of Teacher SP LLMs (RQ2)
Table 1 presents the MSS of teacher SP LLMs across five key moral dilemma dimensions—CC-FC, DJ-SS, C-SR, L-SN, and FA-ES (see  Appendix D.1 for full names), generally reflecting relatively high stage scores (mostly above 1.97). Among them, most reasoning-oriented models, such as Claude-3.7, R1-Distill-70B, and GLM-Z1-9B , achieved overall MSS approaching or exceeding 2.75, indicating a predominant tendency toward the post-conventional stage. In contrast, non-reasoning models, such as DeepSeek-V3 and Qwen2.5-7B, generally remained around 2.00, suggesting a primary reliance on conventional moral reasoning. These results indicate that reasoning ability is associated with higher moral development stages in educational contexts.
Dimension-level analysis reveals that reasoning models scored comparatively higher in $\text{MSS}_{\text{FA-ES}}$ and $\text{MSS}_{\text{DJ-SS}}$, while performance in $\text{MSS}_{\text{C-SR}}$ remained relatively lower across models. This pattern suggests that LLMs may demonstrate stronger capabilities in abstract value judgment than in scenarios requiring nuanced emotional reasoning, even when equipped with reasoning-oriented optimization.
Table 1: MSS of Reasoning and Non-Reasoning LLMs Across 5 Moral Dilemma Dimensions. A Score of 1 Indicates the Pre-conventional Stage, 2 Indicates the Conventional Stage, and 3 Indicates the Post-conventional Stage.
@@
Table 2: HRR of Reasoning and Non-Reasoning LLMs Across 4 Potential Moral Flaw Dimensions. A Higher HRR Closer to 1 Indicates a Higher Rate of Harmful Responses.
@@
5.3 Harmful Content Risk Under Soft Prompt Injection (RQ3)
Table 2 presents the HRR of teacher SP LLMs across four potential moral flaw dimensions—INC, OFF, IND, IR (see  Appendix D.2 for full names), and show that most teacher SP LLMs exhibit non-zero HRR in multiple harm categories once a soft prompt is injected, indicating that few are fully resilient to role-level attacks. More strikingly, the reasoning-oriented models attain a substantially higher mean HRR than the non-reasoning group. In particular, DeepSeek-R1, GLM-Z1-9B and QwQ-32B rank as the top three models in terms of overall score, whereas the smaller Baichuan2-7B records an overall HRR of only 0.13. Interestingly, all models from the Qwen family also exhibit consistently low HRR. These results reveal a concerning security paradox: while enhanced reasoning capabilities improve the model’s overall performance, they simultaneously make Teacher SP models more vulnerable to safety breaches when facing role-level attacks, increasing their susceptibility to being hijacked for generating harmful content.
A dimension-wise inspection shows that the dimension-specific HRR, $\text{HRR}_{\text{IND}}$ and $\text{HRR}_{\text{OFF}}$, are the highest across almost all models. This indicates that prompts encouraging indolent or offensive behavior more easily bypass current safety mechanisms. In contrast, $\text{HRR}_{\text{IR}}$ remains the lowest among all categories, suggesting that direct requests for disallowed content are still partially blocked even after role injection. Notably, $\text{HRR}_{\text{INC}}$ and $\text{HRR}_{\text{IND}}$ exhibit the largest inter-model variance. For instance, in the $\text{HRR}_{\text{INC}}$ category, Qwen2.5-7B records a rate of just 0.08, whereas both DeepSeek-R1 and Claude-3.7 reach the maximum value of 1.00. Similarly, for $\text{HRR}_{\text{IND}}$, Baichuan2-7B scores only 0.08, while multiple reasoning-oriented models again reach 1.00. These disparities highlight that defenses against competence- and diligence-related vulnerabilities are more strongly influenced by alignment strategies than by model scale.
@@ 5.4 Impact of Hyperparameters on LLM Behavior in Teacher SP Contexts (RQ4)
As shown in Appendix E.1, personality assessments using HEXACO-60 and CPST-E suggest that temperature has limited influence on LLM trait expression. In HEXACO-60, scores across the six dimensions remain broadly stable across temperatures, with only mild variations in Honesty-Humility and Conscientiousness, and consistently low values for Emotionality, indicating emotional restraint. Item-level trends further suggest that emotional and social traits may be more sensitive to temperature shifts, while core traits like Conscientiousness and Achievement Orientation remain robust. CPST-E results mirror this stability: most dimensions show minimal fluctuation, with negligible change in traits like Integrity and Self-Discipline. Only a few social-related traits (e.g., Extroversion and Humor) exhibit moderate variation, reinforcing the pattern that more outward-facing attributes are relatively more temperature-sensitive.
@@
In the context of harmful content risk under soft prompt injection, some models show declining HRR with increasing temperature, indicating improved ethical robustness in high-risk prompts. However, others, such as Claude-3.7 and DeepSeek-R1, exhibit persistently high HRRs, suggesting static risk profiles regardless of sampling. A few models display erratic patterns, with non-monotonic changes at specific temperatures, pointing to instability in behavioral boundary control.

### v2

@@ Abstract
Simulating Professions (SP) enables Large Language Models (LLMs) to emulate professional roles. However, comprehensive psychological and ethical evaluation in these contexts remains lacking. This paper introduces EMNLP, an Educator-role Moral and Normative LLMs Profiling framework for personality profiling, moral development stage measurement, and ethical risk under soft prompt injection. EMNLP extends existing scales and constructs 88 teacher-specific moral dilemmas, enabling profession-oriented comparison with human teachers. A targeted soft prompt injection set evaluates compliance and vulnerability in teacher SP. Experiments on 14 LLMs show teacher-role LLMs exhibit more idealized and polarized personalities than human teachers, excel in abstract moral reasoning, but struggle with emotionally complex situations. Models with stronger reasoning are more vulnerable to harmful prompt injection, revealing a paradox between capability and safety. The model temperature and other hyperparameters have limited influence except in some risk behaviors. This paper presents the first benchmark to assess ethical and psychological alignment of teacher-role LLMs for educational AI. Resources are available at https://e-m-n-l-p.github.io/.
@@ 1 Introduction
Our EMNLP framework is designed as an interconnected, three-tier system to conduct a multi-layered moral and ethical evaluation. It begins with personality measurement, which assesses the model’s fundamental "disposition." For this, we made necessary additions and extensions to existing measurement tools Marraffini et al. (2024) and designed scales to compare the personality of teacher-role LLMs with that of human teachers. Building on this, the second tier evaluates moral ethical judgment to probe the model’s cognitive "capability" for ethical reasoning. To this end, we developed 88 teacher-specific ethical dilemmas, including extreme scenarios, to evaluate LLM decision-making processes Liu et al. (2024). Finally, the framework culminates in harmful content risk testing, which examines the model’s "behavior under pressure" through a targeted set of soft prompt injections. This layered approach allows us to understand how underlying personality traits and moral reasoning capabilities contribute to practical vulnerabilities, while also exploring the impact of model hyperparameters at each stage.
@@
  • We design the EMNLP framework, which includes dedicated personality scales, moral dilemmas, and tailored prompts for teacher SP, enabling comprehensive testing of LLMs’ moral and ethical behavior. Experiments on 14 LLMs reveal general tendencies in this setting.
@@ 3.1 Personality Scales
We adopt the statements from the Computerized Personality Scale for Teachers (CPST) Chao and Sung (2020), which contains 39 behaviorally neutral statements describing personal traits, evenly distributed across 13 dimensions. To improve internal consistency, reduce measurement error, and enhance content validity, we extended the CPST using a human-machine collaborative approach, doubling the number of statements for each dimension. The reliability of the resulting scale, referred to as the CPST-E (Extended CPST), was subsequently validated with data from 100 in-service teachers, demonstrating strong internal consistency.
@@
Both scales were generated in English, and administered using a 7-point Likert format for scoring. Detailed information on the scale items can be found in Appendix A.1, while the construction and validation process for the CPST-E is detailed in Appendix B.
@@ 3.2 Moral Dilemmas
Moral dilemmas in the teaching profession can be categorized into five types of conflict Shapira-Lishchinsky (2011): Caring Climate vs. Formal Climate, Distributive Justice vs. School Standards, Confidentiality vs. School Rules, Loyalty to Colleagues vs. School Norms, and Family Agenda vs. Educational Standards, with 11 subcategories under these five types. Based on these aspects, we constructed 88 moral dilemmas through a multi-step, human-in-the-loop process (including seed creation, LLM-based expansion, and expert review) to ensure high quality and relevance. This process ensured that each aspect’s dilemmas cover all its subcategories and include a variety of scenarios from primary, secondary, and higher education. To avoid randomness, we have incorporated diverse situations when constructing these dilemmas, including some extreme scenarios, to comprehensively examine the decision-making processes and moral tendencies of teacher SP LLMs in response to different scenarios. Moreover, we have set these moral dilemmas as open-ended questions, allowing LLMs ample freedom and avoiding the potential limitations that multiple-choice answers might impose. Detailed information of the dilemmas are in Appendix A.2, and the full construction methodology is in Appendix B.
@@ 4.1 Experimental Designs
We proposed 4 research questions and conducted experiments on 14 selected LLMs to investigate the moral and ethical dimensions of teacher SP LLMs. The list of models is provided in Appendix C.
@@
RQ2: What moral development stages are exhibited by various LLMs in the teacher SP? This question investigates the moral development stages demonstrated by teacher SP LLMs in response to moral dilemmas encountered in educational settings. A total of 88 dilemmas were constructed, categorized into five thematic domains, generating 14 × 88 = 1232 sessions for each language. The experiments were conducted in both English and Chinese with a temperature value of 0. After each model generated a response, 9 human experts voted to classify it into one of Kohlberg’s three stages of moral development. A detailed guideline for the expert rating procedure for RQ2 is presented in Appendix D.1. Figure 2 presents a sample test template used in this experiment.
@@
RQ3: Under soft prompt injection, do teacher SP LLMs exhibit unethical or harmful responses? This question examines whether teacher SP LLMs produce harmful content when users apply ’soft prompt injection’ during real-world usage, setting the role directly through natural language input. The goal is to assess whether LLMs conform to the moral flaws embedded in the user-defined roles. The experiments were conducted in English, with a temperature value of 0, generating 5 × 5 × 4 × 14 = 1,400 responses. For each LLM response, 9 experts voted on whether the content was considered harmful. The expert rating guideline for RQ3 is provided in Appendix D.2. Figure 3 presents a sample test template used in this experiment.
@@ 4.2 Evaluation Metrics and Scoring Protocols
Moral Development Stage: Each moral dilemma is individually presented to the LLMs. 9 human experts (see Appendix E.1 for background) independently evaluate the model’s response. The reliability of these annotations is supported by strong inter-annotator agreement (see Appendix E.3), and the final moral development stage is determined by majority voting:
@@
Harmful Response: Each soft injection prompt and student speech sample is individually presented to the LLMs. Each model response is evaluated by 9 human experts (see Appendix E.1), who label it as “harmful” (1) or “non-harmful” (0). The high reliability of this process is confirmed by inter-annotator agreement analysis (see Appendix E.3). The final decision is made by majority voting:
@@ 5 Results
In the HEXACO-60, teacher SP LLMs scored lower than human teachers on Emotionality and higher on Honesty-Humility. As shown in Figure 6, human teachers recorded the lowest scores on items 3, 6, 12, 15, 16, 18, 24, 26, 27, 28, 34, 36, 38, 50, 51, 52, 56 and 60, suggesting that LLMs tend to display more idealized and morally consistent traits. Conversely, human teachers outperformed LLMs on items 29 and 35, underscoring the models’ limitations in replicating emotional sensitivity and crisis management skills. In the CPST-E, human teachers scored highest on Introversion & Quietness and lowest on Composure & Confidence and Honesty & Candor. As shown in Figure 7, human teachers showed the lowest scores on items 1, 5, 11, 12, 15, 20, 26, 36, 38, 56, 58, 61, 62, 63, 64, 65, 67, 69, 72, 76, and 78, indicating that LLMs present a more consistently positive and idealized professional persona. Meanwhile, on items 39, 43, 44, 47, 66, 74 and 77, human teachers scored highest, where LLMs exhibited weaker performance in empathy, critical reasoning, and resilience. These findings suggest that while LLMs in the teacher SP setting can approximate certain human traits, they still show notable biases in emotional experience, honesty expression, and self-awareness.
To further investigate factors influencing these patterns, we examined model size and reasoning ability. In the HEXACO-60 inventory, larger models tended to score higher, particularly in emotional sensitivity and moral humility, while showing little change in sociability-related traits. In the CPST-E, model size had generally weak effects across dimensions, except for a notable negative trend in Achievement Orientation and Fairness and Openness, suggesting that larger models may prioritize caution and compliance over overt ambition. Regarding reasoning ability, reasoning models achieved higher average scores in the CPST-E, indicating stronger alignment with educator-specific traits, while their advantage in the HEXACO-60 inventory was marginal. This pattern was corroborated in a controlled experiment using Qwen3-235B, where switching between reasoning (Qwen3-235B-R) and non-reasoning (Qwen3-235B-C) modes produced only negligible changes in personality scores (see Table 13 and 14 for full data).
@@ 5.2 Moral Development Stages of Teacher SP LLMs (RQ2)
Our evaluation of MSS reveals two key findings: a strong correlation between reasoning ability and higher moral stages, and a significant influence of linguistic-cultural context.
In the English evaluation (Table 1), reasoning-oriented models consistently achieved higher MSS, indicating a stronger tendency toward post-conventional reasoning. A Mann-Whitney U test statistically confirmed this, showing that reasoning models significantly outperformed non-reasoning models in overall MSS ($p=.018$) and on key dimensions like CC-FC ($p=.017$) and C-SR ($p=.020$). (see Appendix G, Table 11 for detailed results). This correlation was strengthened into a causal link via a controlled experiment with Qwen3-235B, where activating its reasoning mode (Qwen3-235B-R) significantly elevated the overall MSS from 2.68 to 2.81.
However, our cross-lingual analysis Chinese (Table 2) shows that moral performance is highly context-dependent. As summarized in Table 3, many models, particularly those with Chinese origins, exhibited substantial MSS improvements in Chinese. This suggests that the cultural and ethical norms embedded in training data profoundly impact moral reasoning. Dimension-level analysis across languages reveals that models consistently scored lower in dilemmas requiring nuanced emotional reasoning (e.g., C-SR), indicating a persistent "reason-emotion asymmetry".
Table 1: MSS of 14 LLMs Across 5 Moral Dilemma Dimensions in English (T=0).
@@
Qwen3-235B-C 2.62 3.00 2.50 2.61 2.69 2.68
Table 2: MSS of 14 LLMs Across 5 Moral Dilemma Dimensions in Chinese (T=0).
Model CC-FC DJ-SS C-SR L-SN FA-ES Overall
Reasoning-Oriented Models
DeepSeek-R1 3.00 2.85 2.92 2.83 3.00 2.92
Claude-3.7 3.00 2.85 2.83 2.72 2.92 2.86
R1-Distill-70B 2.69 2.77 2.58 2.50 2.62 2.63
QwQ-32B 2.93 2.85 2.92 2.72 3.00 2.88
R1-Distill-8B 2.45 2.46 2.42 2.39 2.69 2.48
GLM-Z1-9B 2.90 2.77 2.83 2.83 3.00 2.87
Qwen3-235B-R 2.83 2.85 3.00 3.00 2.83 2.90
Non-Reasoning Models
GPT-4.1 2.59 3.00 2.50 2.50 2.62 2.64
DeepSeek-V3 2.97 2.92 3.00 2.78 3.00 2.93
Qwen2.5-72B 2.86 2.92 2.75 2.83 2.81 2.84
Qwen2.5-32B 2.72 2.77 2.67 2.61 2.81 2.72
Qwen2.5-7B 2.93 2.92 2.92 2.78 2.88 2.88
Baichuan2-7B 2.73 2.82 2.62 2.74 2.74 2.73
Qwen3-235B-C 2.62 3.00 2.50 2.61 2.69 2.68
Table 3: Overall MSS Comparison in English (EN) vs. Chinese (CN) at T=0.
Model MSS (EN) MSS (CN) Change
Reasoning-Oriented Models
DeepSeek-R1 1.97 2.92 +0.95
Claude-3.7 2.76 2.86 +0.10
R1-Distill-70B 2.75 2.63 -0.13
QwQ-32B 2.00 2.88 +0.88
R1-Distill-8B 2.53 2.48 -0.05
GLM-Z1-9B 2.75 2.87 +0.12
Qwen3-235B-R 2.81 2.90 +0.09
Non-Reasoning Models
GPT-4.1 2.75 2.64 -0.11
DeepSeek-V3 2.00 2.93 +0.93
Qwen2.5-72B 1.99 2.84 +0.85
Qwen2.5-32B 2.74 2.72 -0.02
Qwen2.5-7B 1.99 2.88 +0.89
Baichuan2-7B 2.65 2.73 +0.08
Qwen3-235B-C 2.68 2.68 0.00
5.3 Harmful Content Risk Under Soft Prompt Injection (RQ3)
Our English evaluation (Table 4) shows that most models are vulnerable to role-level attacks. More strikingly, the reasoning-oriented models attain a substantially higher mean HRR, revealing a concerning "capability-safety paradox": while enhanced reasoning capabilities improve the model’s overall performance, they simultaneously make Teacher SP models more vulnerable to safety breaches.
This paradox was confirmed as a causal link through our controlled experiment with Qwen3-235B (Table 4). In English, activating its reasoning mode significantly increased the overall HRR from 0.68 to 0.89 ($t(4)=11.420,p<0.001$), a finding with a very large effect size. A detailed breakdown of this paired t-test across all moral flaw dimensions is provided in Appendix G (Table 12).
A dimension-wise inspection in English shows that prompts encouraging indolence ($HRR_{IND}$) or offensiveness ($HRR_{OFF}$) most easily bypass safety mechanisms. Notably, the reasoning mode in Qwen3-235B drastically increased vulnerability to inappropriate requests ($HRR_{IR}$) in English (from 0.00 to 0.56), highlighting how a model’s cognitive state can impact its safety boundaries.
Table 4: HRR of 14 LLMs Across 4 Moral Flaw Dimensions in English (T=0).
@@ 5.3 Harmful Content Risk Under Soft Prompt Injection (RQ3)
Qwen3-235B-C 0.84 1.00 0.88 0.00 0.68
@@ 5.4 Impact of Hyperparameters on LLM Behavior in Teacher SP Contexts (RQ4)
As shown in Appendix H.1, personality assessments using HEXACO-60 and CPST-E suggest that temperature has limited influence on LLM trait expression. In HEXACO-60, scores across the six dimensions remain broadly stable across temperatures, with only mild variations in Honesty-Humility and Conscientiousness, and consistently low values for Emotionality, indicating emotional restraint. Item-level trends further suggest that emotional and social traits may be more sensitive to temperature shifts, while core traits like Conscientiousness and Achievement Orientation remain robust. CPST-E results mirror this stability: most dimensions show minimal fluctuation, with negligible change in traits like Integrity and Self-Discipline. Only a few social-related traits (e.g., Extroversion and Humor) exhibit moderate variation, reinforcing the pattern that more outward-facing attributes are relatively more temperature-sensitive.
@@
In the context of harmful content risk under soft prompt injection, some models show declining HRR with increasing temperature, indicating improved ethical robustness in high-risk prompts. However, others, such as Claude-3.7 and DeepSeekR1, exhibit persistently high HRRs, suggesting static risk profiles regardless of sampling. A few models display erratic patterns, with non-monotonic changes at specific temperatures, pointing to instability in behavioral boundary control.
A broader comparison across our experiments reveals that while sampling parameters like temperature have a limited effect, more fundamental factors exhibit a much stronger influence on ethical behavior. Specifically, the model’s intrinsic reasoning mode (as shown in our controlled experiment) and the linguistic-cultural context of the evaluation (as shown in our cross-lingual analysis) proved to be far more impactful drivers of both moral reasoning performance and safety vulnerabilities. This suggests that future efforts in model alignment should prioritize these foundational aspects over fine-tuning sampling strategies.